# New versions of Frobenius and integral closure of ideals

### v2

@@ 3.4. Comparing the new closures
We claim that $x\in\overline{(I^{d})}^{\mathfrak{a}}.$ Let $\mathfrak{a}=(v_{1},\ldots,v_{\mu}).$ For $1\leq i\leq\mu,$ from (3.1) we know that there exists $c_{i}\in R^{\circ}$ such that $c_{i}x^{n}v_{i}^{n}\in I^{nd}$ for all $n.$ Then for $n_{1}+\cdots+n_{\mu}=n,$
@@
$\left(\prod_{i=1}^{\mu}c_{i}\right)x^{n}v_{1}^{n_{1}}\cdots v_{\mu}^{n_{\mu}}=\prod_{i=1}^{\mu}\big{(}c_{i}x^{n_{i}}v_{i}^{n_{i}}\big{)}\subset I^{nd},$
@@
implying that $x\in\overline{(I^{d})}^{\mathfrak{a}}\subset I^{*\mathfrak{a}}$ using Theorem 3.19.
@@
Now suppose that not all $f_{i}$’s have the same degree. Choose nonnegative integers $\alpha_{i}$ such that $F_{1}=f_{1}^{\alpha_{1}+1},\ldots,F_{d}=f_{d}^{\alpha_{d}+1}$ have equal degrees, say $D$. Let $x\in R_{N}$ where $N\geq\sum_{i=1}^{d}\deg f_{i}-\beta^{\prime}l.$ Then $xf_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}}$ has degree at least $dD-\beta^{\prime}l$ and hence by the first case is in $(F_{1},\ldots,F_{d})^{*\mathfrak{a}}.$ By definition, there exists $c\in R^{\circ}$ such that for all $e$ large,
$c(xf_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}})^{p^{e}}\mathfrak{a}^{p^{e}}\subset(f_{1}^{(\alpha_{1}+1)p^{e}},\ldots,f_{d}^{(\alpha_{d}+1)p^{e}}).$
@@
$cx^{p^{e}}\mathfrak{a}^{p^{e}}\subset(f_{1}^{(\alpha_{1}+1)p^{e}},\ldots,f_{d}^{(\alpha_{d}+1)p^{e}}):(f_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}})^{p^{e}}\subset(f_{1}^{p^{e}},\ldots,f_{d}^{p^{e}})^{*}$

### v3

@@ 4.3. Adjusted closures in graded rings
We claim that $x\in\overline{(I^{d})}^{\mathfrak{a}^{t}}.$ Let $\mathfrak{a}=(v_{1},\ldots,v_{\mu}).$ For $1\leq i\leq\mu,$ we know that there exists $c_{i}\in R^{\circ}$ and $r_{i}\in\mathbb{N}$ such that $c_{i}(x^{r_{i}}v^{\lfloor r_{i}t\rfloor})^{n}\in I^{nr_{i}d}$ for all $n$. Fix $n_{1},\ldots,n_{\mu}\in\mathbb{N}$ such that $n_{1}+\cdots+n_{\mu}=\lceil tn\rceil,$ i.e. $v_{1}^{n_{1}}\cdots v_{\mu}^{n_{\mu}}$ is a generator of $\mathfrak{a}^{\lceil tn\rceil}$. For each $i=1,\ldots,\mu$, let $q_{i}=\lfloor n\cdot n_{i}/\lceil tn\rceil\rfloor$. Note that $\sum_{i}q_{i}\leq\sum_{i}n\cdot n_{i}/\lceil tn\rceil=n$. Also, note that $n_{i}/(r_{i}t)=nn_{i}/(tnr_{i})\geq nn_{i}/(\lceil tn\rceil r_{i})\geq q_{i}/r_{i}$.
Now we have that for any $i$,
$c_{i}x^{q_{i}}v_{i}^{n_{i}}\in c_{i}\left(x^{r_{i}\lfloor q_{i}/r_{i}\rfloor}v_{i}^{\lfloor r_{i}t\rfloor\lfloor n_{i}/(r_{i}t)\rfloor}\right)\subset c_{i}\left(x^{r_{i}\lfloor q_{i}/r_{i}\rfloor}v_{i}^{\lfloor r_{i}t\rfloor\lfloor q_{i}/r_{i}\rfloor}\right)\subset c_{i}\left(x^{r_{i}}v_{i}^{\lfloor r_{i}t\rfloor}\right)^{\lfloor q_{i}/r_{i}\rfloor}\subset I^{\lfloor q_{i}/r_{i}\rfloor r_{i}d}.$
Therefore,
$\left(\prod_{i=1}^{\mu}c_{i}\right)x^{n}v_{1}^{n_{1}}\cdots v_{\mu}^{n_{\mu}}\in\prod_{i=1}^{\mu}\big(c_{i}x^{q_{i}}v_{i}^{n_{i}}\big)\subset\prod_{i=1}^{\mu}I^{\lfloor q_{i}/r_{i}\rfloor r_{i}d}=I^{d\sum_{i}\lfloor q_{i}/r_{i}\rfloor r_{i}}.$
Now
$\sum_{i=1}^{\mu}\left\lfloor\dfrac{q_{i}}{r_{i}}\right\rfloor r_{i}>\sum_{i=1}^{\mu}\left(\dfrac{1}{r_{i}}\left\lfloor\dfrac{n\cdot n_{i}}{\lceil tn\rceil}\right\rfloor-1\right)r_{i}>\sum_{i=1}^{\mu}\left(\dfrac{n\cdot n_{i}}{\lceil tn\rceil}-1-r_{i}\right)=n-\sum_{i=1}^{\mu}(r_{i}+1),$
hence $d\sum_{i}\lfloor q_{i}/r_{i}\rfloor r_{i}>nd-d\sum_{i}(r_{i}+1)$. Therefore, taking $0\neq c^{\prime}\in I^{d\sum_{i}(r_{i}+1)}$, we have that for all $n\geq 1$,
@@
$c^{\prime}\left(\prod_{i=1}^{\mu}c_{i}\right)x^{n}\mathfrak{a}^{\lceil tn\rceil}\subset I^{nd}$
@@
implying that $x\in\overline{(I^{d})}^{\mathfrak{a}^{t}}\subset I^{*\mathfrak{a}^{t}}$ using Theorem 4.5. Note that for this case we do not require the existence of a parameter test element or equidimensionality.
@@
Now suppose that not all $f_{i}$’s have the same degree. Choose nonnegative integers $\alpha_{i}$ such that $F_{1}=f_{1}^{\alpha_{1}+1},\ldots,F_{d}=f_{d}^{\alpha_{d}+1}$ have equal degree $D$. Let $x\in R_{\geq N}$. Then $xf_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}}$ has degree at least
$\sum_{i=1}^{d}\deg f_{i}-t\beta^{\prime}l+\sum_{i=1}^{d}{\alpha_{i}\deg(f_{i})}=dD-t\beta^{\prime}l.$
Hence, by the first case, $xf_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}}\in(F_{1},\ldots,F_{d})^{*\mathfrak{a}^{t}}.$ By definition, there exists $c\in R^{\circ}$ such that for all $e$ large,
$c(xf_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}})^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(f_{1}^{(\alpha_{1}+1)p^{e}},\ldots,f_{d}^{(\alpha_{d}+1)p^{e}}).$
@@
$cx^{p^{e}}\mathfrak{a}^{\lceil tp^{e}\rceil}\subset(f_{1}^{(\alpha_{1}+1)p^{e}},\ldots,f_{d}^{(\alpha_{d}+1)p^{e}}):(f_{1}^{\alpha_{1}}\cdots f_{d}^{\alpha_{d}})^{p^{e}}\subset(f_{1}^{p^{e}},\ldots,f_{d}^{p^{e}})^{*}$